OpenStudy (anonymous):

A charge of 6.90 pC is spread uniformly throughout the volume of a sphere of radius r = 4.14 cm. What is the magnitude (in N/C) of the electric field at a radial distance of (a) 6.16 cm and (b) 3.19 cm?

OpenStudy (anonymous):

You can find it via gauss law\[E \ dA =Q \div \epsilon \]

OpenStudy (anonymous):

where \[dA = 2\times \pi \times r\]